The cimento L^{A}TEX class for Il Nuovo Cimento^{∗}

A SIF production written by S. Frabetti

See sec. 5 for addresses.

Printed January 11, 2002

Abstract

This document describes the L^{A}TEX class provided by the Italian Physical Society (SIF) for typesetting papers for its journal “Il Nuovo Cimento”.

Contents

1 Foreword 2 1.1 How to obtain this file. . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Distribution files and installation . . . . . . . . . . . . . . . . . . . 2

2 Desperate introduction 3

3 Class description 4 3.1 Author vs. editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2 The preamble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3.2.1 Preamble scheme . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2.2 Preamble macros . . . . . . . . . . . . . . . . . . . . . . . . 5

3.3 The main text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.4 Final stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.5 Additional macros . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.6 Things of no consequence . . . . . . . . . . . . . . . . . . . . . . . 13

4 More things of consequence 13 4.1 Figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.2 Style guide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.3 Varenna . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.4 Rivista del Nuovo Cimento . . . . . . . . . . . . . . . . . . . . . . 15

5 Contacting 15

6 A sample paper 15

∗This file has version number v1.4c, last revised 1999/12/01.

1 Foreword

The class cimento is a variation of the standard L^{A}TEX article class, especially designed to produce an output that matches the standards of the review “Il Nuovo Cimento”.

Papers written in L^{A}TEX and based on this class will undergo the publication process more easily, more quickly and with less structural changes than in any other possible format, such as plain TEX, Word, WordPerfect, etc.

While the printed appearance of a paper based on this class will be quite different from that of a standard L^{A}TEX article, we have tried to preserve as far as possible the syntax of L^{A}TEX, so that you won’t have to learn many new commands and only minor changes should be necessary to adapt a text already written in L^{A}TEX.

Please notice that this class is still under development and it changes often: you should try to get the latest version before starting to write a paper based upon it.

1.1 How to obtain this file.

This package is on line; get it by FTP at the address

     ftp.sif.it/pub/tex/cimento/

and on the web at

     www.sif.it.

You can also have it on a disquette by making a request at one of the addresses listed in section 5.

Finally, this package is free: use it, distribute it, even modify it if you feel like—we cannot prevent it. There is a copyright notice in the distribution file

cimento.dtx.

1.2 Distribution files and installation

The package is distributed as file \underline{cimento.dtx}; it is a single docstrip file, that is, a L^{A}TEX file which contains both the source code for the class and its documentation (i.e., what you are reading). To obtain this documentation, you just have to compile cimento.dtx using L^{A}TEX as you would with any other L^{A}TEX source file—just do it thrice, to be sure that the page numbers in the table of contents are right.

On the other hand, to obtain the actual class file to be called by your own L^{A}TEX source, you have to strip it from the dtx file; you can do this with the docstrip program, if you know how to use it (options class and sample), but the easiest way to do this is by using the installation file included in the distribution, \underline{cimento.ins}: type latex cimento.ins. That will extract two files from

cimento.dtx: \underline{cimento.cls} and \underline{cimsmple.tex}. The first contains the code for

the cimento class; this is all you really need to prepare your paper. Since L^{A}TEX loads this file whenever you compile your paper, it must reside in a directory searched by the program, so you have to put it either in the L^{A}TEX input directory, or in the directory where the file you are actually writing dwells^{1}.

The second file extracted from cimento.dtx, cimsmple.tex, is a short sample paper that you can compile to compare the output with the source code; you could also make a copy of it and use it as a template for your own paper. Section 6 contains a commented copy of the source code of this sample.

The following table resumes the filenames and their contents

distribution files

README.TXT  readme file…

cimento.dtx the “compressed” file which contains the class

code along with the document you are reading

cimento.ins  a script file for unpacking the distribution

unpacked files

cimento.cls the actual class file, the one you will input in your

paper

cimsmple.tex  a sample/template paper

2 Desperate introduction

If you can tell the difference between L^{A}TEX 2.09 and L^{A}TEX 2ε please skip this section; otherwise, this is as good a place as any other to learn something about that.

L^{A}TEX 2ε is the new version of L^{A}TEX which has been released in 1994. Since it has replaced the older version, it is of course refered to simply as ‘L^{A}TEX’, and so is in this document. Whenever there is the necessity to distinguish this new version from the older one, the latter is called ‘L^{A}TEX 2.09’, from its version number.

Needless to say, the cimento package is based upon the new version; you should seriously consider the opportunity of documenting yourself about the features of this new version of L^{A}TEX. See L^{A}TEX: A Document Preparation System [1] or The L^{A}TEX companion [2] for a thorough treatment of the subject.

However, some facts need to be pointed out here:

• L^{A}TEX 2ε behaves quite well with respect to documents written for L^{A}TEX 2.09, but,

• in order to distinguish clearly between old macro packages and new ones, the former keep their traditional name of “style files” and the .sty extension, while the latter are called “class files” with extension .cls—this explains why we call cimento a “class” instead of a “style” file.

• The first line of a document makes L^{A}TEX decide whether it’s facing a new document or an old one:

\documentstyle is L^{A}TEX 2.09 incipit; \documentclass is L^{A}TEX 2ε new one.

• Whenever you write in L^{A}TEX, you should try and start your paper with

1Note that the former choice may not be available to you: if you work in a Unix environment, you probably don’t have write access rights to the L^{A}TEX directory. You can still play with environment variables to have your TEX search in your own user-defined path…

         \documentclass{article};

this will trigger L^{A}TEX 2ε mode, which is more convenient—much faster, for instance. Then, if you don’t want to spend some time learning the new wonderful commands at your disposal, you can just get along with the commands you are accustomed to; the minor syntax differences are conveniently handled for you by L^{A}TEX itself.

Now, this package is written for L^{A}TEX 2ε; that means you must begin your paper with the new starting command, like this:

     \documentclass{cimento}.

To conclude this section, we point out that L^{A}TEX 2ε is the standard now, and as such is probably already installed on your system (if not, complain!); that’s why you had better cope with it and take a glance at one of the books cited above—look at your local library first.

3 Class description

The class is really a standard article, with but a little make-up; therefore the description which follows is simply a list of the peculiarities and differences between this class and the standard one.

3.1 Author vs. editor

The cimento class can work in two different major modes: draft mode and final mode. The first one is used to produce a draft of the paper for first reading, while the latter produces an output (of the same paper!) which matches closely the printed appearance of the review. You can choose the mode that best fits you to prepare your paper; all we’ll have to do is switch to the final mode when the file arrives to the editorial office.

In any case, we recommend that you don’t misuse the final mode by getting involved with visual formatting, like inserting \vspaces, \pagebreaks, \linebreaks etc. This is very important; for example, every line break that you put in to achieve a better visual appearance will be deleted as soon as it is found, and this is true for the greatest part of the low-level formatting macros. (As another example, there are papers with the math cluttered up with \!s, \,s, etc.; this is definitevely bad!)

Moreover, stick to L^{A}TEX whenever possible; as far as you can, shun plain TEX. Up to now, we are not sure of the usefulness of the draft mode: the default

is final and you have to explicitely request the draft if you want it (see below for how to do this).

3.2 The preamble

We first give a synoptic scheme of a typical preamble, then we will proceed to describe each macro individually.

3.2.1 Preamble scheme

     \documentclass[options]{cimento}
     \usepackage{package1}
     \usepackage{package2}
     ...
     \newcommand{comm1}{...}
     \newcommand{comm2}{...}
     ...
     \title{paper title}
     \author{first author\from{i1},
             second author\from{i2},
             ...
             \atque last author\from{in}}
     \instlist{\inst{i1} first institute
               \inst{i2} second institute
               ...
               \isnt{in} last institute}
     \PACSes{\PACS{first ID}{first description}
             \PACS{second ID}{second description}
             ...
             \PACS{last ID}{last description}}
     \begin{document}

3.2.2 Preamble macros

Class options. As stated above, you should start your document with

\documentclass        \documentclass[⟨options⟩]{cimento}.

The options supported by the class are summarized in the table below. All options not explicitly mentioned are not supported and produce a warning message. However, they are kept in memory for packages that may use them, so that, for example, you can still declare an option dvips in the call to \documentclass and have it passed on to the packages.

option  description

author options draft Draft mode: 12pt fonts, stretched page layout, no

running heads and so on.

final The counterpart of the preceding one. It typesets the paper with the final formatting. This one is on by default.

seceqno Numbers equations within sections, rather than

within all document.

cite Load the cite package. bareref Alternate mode for references. Not documented

still!

varenna Varenna mode, see sec. 4.3 below. rivista Rivista mode, see sec. 4.4 below. editorial options overfull Gives overfull marks on the right margin, as in

              draft mode.

shortnote Used to produce a short note. Like the next two,

this is still quite experimental.

erratum Used to produce errata. toc Table of contents of an issue (formerly index). prefatory Prefatory matter: introductions, prefaces, etc. century This option substitutes the standard TEX fonts

with postscript fonts. You don’t have these fonts, so don’t bother yourself with trying it out.

debug Makes TEX error messages more detailed. Useful mainly while debugging the class code, i.e., not useful for you!

The only default option is final.

As you can see, the class options are generally not very useful to authors; you can safely ignore them and begin your paper simply with

\documentclass{cimento}.

Title and author. The \title^{2} macro in draft mode is apparently the same as in article, but since it does more than you can tell at first glance, you have to use it, together with \maketitle, and do not build the title up by means of low-level formatting commands.

The same is true for \author.

\atque Inside the \author command use \atque to separate the last author in a list;

this will produce an ‘and’ that will not get uppercased in the running head.

\thanks To place a footnote in either the title or the author list use \thanks like in standard L^{A}TEX; don’t forget to \protect fragile commands in its argument.

Running heads. The running heads do not appear in draft mode, so if you do not use the final mode you can happily skip this section and proceed to the description of the institutes below.

2It is merely a matter of taste whether these commands (and the subsequent ones in this section) should reside in the preamble or in the main text, after the \begin{document}; the point is that they should go before \maketitle.

The \title macro in fact provides both the main title of the paper and the short title for the running heads. If the title is too long to fit into the running

\ETC head, a \ETC macro has to be inserted inside the title at the point where it should be cut. For example,

     \title{All you ever wanted to know\ETC\ about collapsing
            meteors, but were afraid to ask}

would produce

All you ever wanted to know about collapsing meteors, but were afraid to ask

as a title, but just

ALL YOU EVER WANTED TO KNOW ETC.

in the running head. A backslash after \ETC is necessary if a space has to appear at that spot^{3}. \ETC ignores preceding blanks and provides automatically the necessary spacing.

With respect to the running head and the \ETC macro, \author behaves exactly the same way as \title. Note that its use is required only when the authors are too many to fit into the running head, e.g. not for two authors!

When the short title for the running head should not be obtained simply by cutting the main title, but should instead be a (partly or completely) different text,

\shorttitle you should ignore the \ETC mechanism, and use, instead, \shorttitle{⟨running-head-title⟩}.

   Do not use \ETC inside \shorttitle.

\shortauthor \shortauthor{⟨running-head-author⟩} does the same job for the author, although its use should be quite rare. Do not use \ETC inside \shortauthor, either.

Important note. The text of the title (and author) gets capitalized in the running heads. All the text, even the symbols in mathematics so that, for example,

     \title{The spirit of $E=mc^2$}

becomes

THE SPIRIT OF E = MC^{2}

in the running heads. To avoid this you have to enclose all the letters of the title which should remain lowercase into a \MakeLowercase, like this:

     \title{The spirit of $E=\MakeLowercase{mc}^2$}

We know that this is boring, but we couldn’t see a better solution. (Comments and suggestions are welcome.)

Institutes. To mark an author’s affiliation to an institute, a mechanism is provided which is similar to the \bibitem/\cite one: you declare a list of institutes

\instlist using \instlist{⟨insts⟩}, where ⟨insts⟩ is of the form ⟨insts⟩ := \inst{⟨label-1 ⟩} ⟨institute-description-1 ⟩

3But a comma at that spot should not go into the running head; cfr. this: \title{On collapsing meteors\ETC, blind asteroids and satellites of love.}

\inst{⟨label-2 ⟩} ⟨institute-description-2 ⟩ … \inst{⟨label-n⟩} ⟨institute-description-n⟩

\inst The \insts print the ⟨institute-descriptions⟩ and store away the ⟨labels⟩, which

are then (or rather, were) used to reference the institutes from inside the \author \from argument by means of \from tags: for example,

     \author{A.~Bab\`a\from{ins:gen}} \atque Mr. Albatros\from{ins:air}
     \instlist{\inst{ins:gen} Genius Lamp,
                              1001 Sesamo str. - The Cave
               \inst{ins:air} Main Stream - Troposphere}

will produce something like

A. Babà(^{1}) and Mr. Albatros(^{2})

(^{1}) Genius Lamp, 1001 Sesamo str. - The Cave

(^{2}) Main Stream - Troposphere

The \instlist may appear anywhere before \maketitle (it usually goes after the \author); anyway, it will be printed right below the authors.

Actually, the \instlist construction admits a second syntax, to be used when there is no need of marks to refer to institutes (this is true, for example, when there is but one author, or when all the authors are affiliated to the same institute(s)).

In practice, this secondary syntax is characterized by the absence of \froms in the \author declaration. When this is the case, you just have to omit the ⟨label⟩ argument from all the \insts, so that the syntax of the argument for \instlist becomes

⟨insts⟩ := \inst ⟨institute-description-1 ⟩ \inst ⟨institute-description-2 ⟩ … \inst ⟨institute-description-n⟩

\PACSes Physics and Astronomy Classification Scheme. The \PACSes command

(one argument) holds the PACS list. Each PACS item should reside in a separate

\PACSit \PACSit, whose syntax is:

\PACSit{⟨PACS-ID⟩}{⟨description⟩}

The PACS will be stored and typeset at the end of the abstract. (So put them anywhere before the abstract.)

3.3 The main text

abstract Sectioning. The abstract environment has nothing peculiar (apart from its printed appearance, of course). Sometimes the abstract is too large to fit in the first page; in such cases, it can be split between the first and second page with a

\abstractbreak  \abstractbreak .
\section  The sectioning commands are \section, \subsection, \subsubsection,

\subsection \paragraph and \appendix; all these commands behave exactly as in standard

\subsubsection
\paragraph
\appendix

L^{A}TEX. However, all these commands are redefined to provide Cimento-like section headers.

Mathematics. seceqno is a class option that makes your equations numbered within the sections (e.g. (1.1), (1.2), …, (2.1), etc.):

      \documentclass[seceqno]{cimento}

eqnletter There is also a new environment eqnletter which produces “number-lettered” systems of equations, like this:

\begin{aligned}(1a) & x & = & y \\ (1b) & y & = & z \\ (1c) & z & = & x \end{aligned}

The code for the preceding system is:

     \begin{eqnletter}
          \label{eq.sys}
          x & = & y  \label{eq.sys.r1} \\
          y & = & z  \label{eq.sys.r2} \\
          z & = & x  \label{eq.sys.r3}
     \end{eqnletter}

Note that the first label (positioned before or “outside” of the rows) generates a reference to the whole system (i.e., (\ref{eq.sys}) gives “(1)”), while the next labels, one for each row, store reference information for the single rows

(i.e., (\ref{eq.sys.r1}), (\ref{eq.sys.r2}), (\ref{eq.sys.r3}) give, re-

spectively, “(1a)”, “(1b)” and “(1c)”). The labels, if present, must be in the right position: the general label right after the \begin{eqnletter}; the row labels in the third column of each row: after the second & and before the \\.

Note. There is a package subeqnarray^{4} (it can be found as a supported contribution to L^{A}TEX in the dedicated subdirectory of CTAN) which provides the same functionality of this environment. If you prefer to use it, you may feel free to do it, but you should redefine the way letters are printed, since we want them italic, while they come out of the box roman. You can do it as follows:

     \usepackage{subeqnarray}
     \renewcommand{\thesubequation}{\theequation\textit{\alph{subequation}}}

Citations. Citations like \cite{b:3-b:6}, which should produce ‘[3-6]’, are not permitted by standard L^{A}TEX; however, there is a package which extends the functionality of the \cite command to permit such citations. The package is named cite.sty^{5} and is available in any CTAN mirror (for example, ftp.dante.de) in the directory

     /tex-archive/macros/latex/contrib/supported/cite/

4By J.L.Braams@research.ptt.nl

5by asnd@reg.triumf.ca

(or, ask me and I will send it to you).

We strongly encourage its use, and there is an option [cite] for the cimento class which automatically loads the package.

If you don’t want to install this package, please follow this rule: Don’t type \cite{a}-cite{f}; type instead \cite{a,b,c,d,e,f}: you will get ‘[1,2,3,4,5,6]’, we will get ‘[1-6]’ and everybody will be happy.

Appendices. There are three possible styles for the appendices:

• plain; only one appendix, not “numbered”, like this:

Appendix

This is obtained with an empty non-numbered section:

                 ...preceding text
                 \appendix
                 \section*{}
                 appendix text follows...

• “numbered”; appendices are “numbered” (with uppercase letters!) like normal sections, but they do not have a text header:

Appendix A.

This is the result of using an empty section:

                 ...preceding text
                 \appendix
                 \section{}
                 appendix text follows...

• full header:

Appendix A.

ASCII character table

This is obtained this way:

                 ...preceding text
                 \appendix
                 \section{ASCII character table}
                 follows the table...

Note that a \section*{} (first case above) does not get added to the table of contents. This must be done manually with \addcontentsline as follows:

  \addcontentsline{toc}{section}
                  {\protect\numberline{\normalfont\textsc{Appendix}}{}}

Note also that papers for Il Nuovo Cimento do not have any table of contents… Note. Previously released versions of this file did not allow unnumbered sections

(\section*, \subsection* and \subsubsection*); now they are fully supported.

Tables. Since tables in Nuovo Cimento are always as large as the page, the

tabular tabular environment is redefined (in terms of tabular*…) to span the whole text width; if a different width is really needed, one has to resort to tabular* (but such a table format will most probably be changed during proof-reading). Note that tabular* is left unchanged.

However, if a table has only two or three columns and these are quite narrow so that the table looks a bit too “spaced”, extra space should be allocated beside the leftmost and rightmost column. Nevertheless, the width of the table in itself should remain as large as the page so that, for example, the horizontal rules span the whole page.

narrowtabular The narrowtabular environment does this. Its syntax is:

     begin{narrowtabular}[⟨pos⟩]{⟨colwidth⟩}{⟨cols⟩}

where ⟨pos⟩ and ⟨cols⟩ are the ordinary tabular arguments and ⟨colwidth⟩ is an extra mandatory argument that specifies the amount of horizontal space to be put at the left of the leftmost column and at the right of the rightmost. ⟨colwidth⟩ should be a TEX length (1cm, 2pc etc.).

Drawbacks. The spreading of the tables is achieved through the @-expression

     @{extracolsep{\hfill};

(if you don;t know anything about @-expressions, please ignore this paragraph). Unfortunately, this means that all @-tricks you should eventually insert would end up by adding stretchable white space before the next column so that, in particular, alignment of numbers to the decimal point with the r@{.}l construct is impossible^{6}.

However, two characters are activated inside the tabular environment which ! should help you overcome any aligning difficulty: ! inserts an \hphantom with ? the width of a number; ?{} builds a phantom as large as what you type as an

argument. (To obtain a literal ‘!’ and a literal ‘?’, type, resp., \fact and \?.)

For those who did not ever care to learn what a phantom is, we give here a practicle example of how to align numbers:

      \begin{tabular}{lcl}
          Spinaci   & 3250?{.}!! & lire\\
          Caramella & !!50?{.}!! & lire\\
          X-Men     & !!!1.25    & dollars\\
      \end{tabular}

I used the preceding code to produce the following:

Spinaci  3250  lire
Caramella  50  lire
X-Men  1.25 dollars

By the way, if you think you could make use of ! and ? in other contexts as

\phantoms well, you can gain access to them by declaring \phantoms in your code. There is no \nophantoms available, so take care to include everything in a group, like this:

{\phantoms ...}.

Finally, please note that

• the tabular as here redefined is compatible with the array package and its relatives, so that, for example, you can use dcolumn to align numbers to the dot;

• as explained above, you can obtain weird effects if you try to put vertical lines, @-expressions or other things between columns; however, no vertical lines are admitted in the review tables, so this should not constitute a problem for you.

3.4 Final stuff

Thanks, acknowledgments and the like go at the end (just before the references);

\acknowledgments to separate them from the main text, put a \acknowledgments in between. Note that this command automatically takes care of ending the preceding paragraph and beginning a new one afterwards, so the number of blank lines you leave around

\stars it is irrelevant. \stars is a synonim for \acknowledgments…

thebibliography The thebibliography environment is unchanged, but since references are exclusively tagged with numbers, like [10], you should stick to digit arguments for this environment, such as {0} {00} {000}. For the same reason, the optional

\bibitem argument of \bibitem is disabled.

To help you type the references in a format as close as possible to the review standard, a handful of macros is provided; of course, you may as well ignore them and directly type the text —with the right formatting, we hope. For an example of well formatted references, refer to the sample paper included in the package.

\BY \BY{⟨author(s)⟩} typesets its argument in small caps and appends a comma— don’t type yourself commas, semicolon and the like after this macro. Don’t type ‘and’ before the last author, use instead the macro \atque which produces a lowercase ‘and’ even in a small caps context. Similarly, \etal expands to ‘et al.’

\IN \IN{⟨review⟩}{⟨volume⟩}{⟨year⟩}{⟨page⟩} is for the common review reference. Here again, don’t use any comma, parentheses or similar stuff to separate the fields: it is automatic! For those who care, the ⟨review⟩ is simply italicized; the ⟨volume⟩ is bold-ized; the ⟨year⟩ is enclosed in parentheses; the ⟨page⟩ is left untouched; finally, the first two arguments are separated by a comma.

\SAME \SAME{⟨volume⟩}{⟨year⟩}{⟨page⟩} is used for multiple references to the same review.

\TITLE \TITLE{⟨book⟩} is a ‘low-level’ macro; it is but a synonim for \textit{}; use it for publications which do not match the review-like specifications.

Everything that does not fit into these schemes will have to be typed literally. Actually, a macro is supplied to deal with the unpredictable: \ADLIB{⟨anything⟩}.

\ADLIB This one does absolutely nothing, so it can be used for everything :–)

6unless you manually turn off the hfilling right before your r@{.}l and turn it on again right after.

3.5 Additional macros

A few macros are defined that could be of some help in writing for Il Nuovo Cimento.

\atque \atque expands to ‘and’ in lowercase normal type, even in different contexts, such as small caps, bold, etc. Particularly useful in author and bibliography input. NB: you don’t need to put a slash-space after this one (like this: \atque\ ); it’s automatic. This is true only for this macro.

\etal In a similar way, \etal and \idest expand, respectively, to ‘et al.’ and ‘i.e.’. \idest Another abbreviation: \drm gives an upright (roman) ‘d’ to be used for differ-

\drm entials: $\drm x$ gives dx. Math mode only. \sy \sy{...} is a shorthand for \mathrm{...}; it is used inside equations to insert

letters which do not represent ‘variables’, but ‘names’ (e.g. r_{m}, m = max, as opposed to r_{i}, the i-th r). Math mode only.

\tx \tx{...} is a shorthand for \mbox{...}; use it to insert words or sentences

inside an equation. Both math and paragraph mode. \un \un{...} is a shorthand for \,\mbox{...}; used to typeset measure units:

120\un{GeV} gives 120 GeV. Both modes.

3.6 Things of no consequence

Editorial macros. Two commands are defined which are for editorial use only: \issue \issue{⟨volume⟩}{⟨issue⟩}{⟨date⟩} is used to typeset the running head of the

\checkin first page; \checkin{⟨rec-date⟩}{⟨appr-date⟩} records the dates of submission \checkrec and of acceptance of the paper. There are also a \checkrec{⟨rec-date⟩} and \checkrev a \checkrev{⟨rec-date⟩}{⟨rev-date⟩}{⟨appr-date⟩}.

quotation Customized spacing. The formatting of the quotation environment is slightly changed to match Nuovo Cimento standards, but this is really none of your business… Moreover, we have reduced the amount of space that is put between

eqnarray columns of an eqnarray or eqnarray* environment. For example, the distance between ‘y’ and ‘=’, or between ‘=’ and ‘x’ in

y = x^{2} + 2x + 1
= (x + 1)^{2}

is reduced as in

y = x^{2} + 2x + 1
= (x + 1)^{2}

\footnotesize Font dimension. The \footnotesize command produces the same font size

\small of \small. The command to obtain the font size assigned by standard L^{A}TEX to

\smaller  \footnotesize is \smaller.

4 More things of consequence

There are a few more things you should be aware of.

4.1 Figures

There is up to now no particular facility for inserting figures in cimento documents; you should instead make use of standard L^{A}TEX packages such as graphics or epsf or epsfig or the like. There are some guidelines you should adhere to to help us grant a good printing quality—see below.

4.2 Style guide

While most of the macros (new ones and modified ones) of the class here described were explicitly designed to help you stick to the review style, this document was not meant to be a style guide and much more could be said about style than could possibly go in here.

In fact, a style guide for Il Nuovo Cimento exists, where such idiosyncrasies as title style, reference style, roman and italic letters in formulas, figure production, etc. are covered in detail. This style guide is currently on www at www.sif.it.

4.3 Varenna

The Italian Phisical Society publishes the volumes of proceedings of the Summer School of Physics “Enrico Fermi” (“Varenna school” for short, from the village where it takes place on the lake of Como).

The layout of the volumes is very similar to that of the journal Il Nuovo Cimento; that’s why a L^{A}TEX class for the volumes is presently implemented as a variation of the cimento class.

There is an option [varenna] which switches the class to the Varenna mode, so that the first line of a Varenna paper becomes

     \documentclass[varenna]{cimento}

Authors and institutes. The only relevant point in which the varenna mode differs from the normal cimento mode is in the author/institute syntax.

In the cimento mode, you first list all the author, then all the institutes, using \from/\inst tags to explicit the affilitations.

In varenna mode instead, you have to write all the couples of authors/correspondinginstitute one after the other, and no \from/\inst mechanism is necessary (nor is it available!); so the preamble of a varenna document should look like this:

    \documentclass[varenna]{cimento}
    \title{Title of the paper}
    \author{A. Author, B. Author \atque C. Author}
    \institute{Institute of the First Three Authors}
    \author{X. Author \atque Y. Author}
    \institute{Institute of the Other Two Authors}
    \begin{document}

There may be as many author/institute couples as necessary. Note that the

\institute macro for institutes is named \institute —not \instlist as in cimento mode.

Other differences.

• The command \ETC does not work when there is more than one \author.

\shortauthor If this is your case, you have to use \shortauthor . Do not use \ETC inside

     \shortauthor.

• There is no abstract in Varenna volumes, so don’t use the abstract

abstract environment.

• Something else… ?

4.4 Rivista del Nuovo Cimento

There is an option to the \documentclass that selects the Rivista mode:

    \documentclass[rivista]{cimento}

This mode is still under development.

5 Contacting

Feedback of any kind is greatly appreciated, indeed. If you have any

• question,

• bug report,

• hint,

• anything else

about this package, or related to Il Nuovo Cimento and L^{A}TEX, please write to

stefano@sif.it.

For info about Il Nuovo Cimento in general, that is, not concerning this package in particular (or L^{A}TEX), please write to cimento@sif.it.

La Società Italiana di Fisica is sif@sif.it.

Finally, our web page is

     www.sif.it.

That’s all.

6 A sample paper

Hereafter is included and commented a sample paper which is specifically meant to stress all the peculiarities of the cimento class.

You can also extract it from the distribution file using the docstrip program with option sample (if you use cimento.ins to unpack the distribution (see sec. 1.2), this is done automatically). Once extracted, you can compile it (latex cimsmple) and compare the output with the source code.

1 ⟨∗sample⟩

 2 \documentclass{cimento}

The title below is short enough to fit into the running head; that’s why no \ETC macro is used in it.

 3 \title{Sample paper for the {\mdseries\ttfamily cimento} class}

The authors, on the contrary, are too many, so just a few (just one, this time) are sent to the running head. Notice the comma after the \ETC in the first line.

 4 \author{S.~Summers\from{ins:x}\ETC,
 5 J.~Grey\from{ins:x},
 6 H.~McCoy\from{ins:x},
 7 R.~Drake\from{ins:x}\\
 8 W.~Worthington\from{ins:x}
 9  \atque
10 Mr.~M\from{ins:evil}\thanks{The bad fellow.}}

Also remark the \froms above, corresponding to \insts below.

11 \instlist{\inst{ins:x} Xavier's school for gifted youngsters - Westchester,
12  N.Y.
13 \inst{ins:evil} The Brotherhood of Evil Mutants\-Russel Square,
14  London, UK}

PACSes are stored and typeset inside the abstract box.

15 \PACSes{\PACSit{00.00}{By the way, which PACS is it, the 00.00? GOK.}
16 \PACSit{---.---}{\ldots}}

Usage of the next commands is absolutely standard. Just notice the absence of punctuation in section and subsection headings.

17 \begin{document}

18

19 \maketitle

20

21 \begin{abstract}
22 This sample paper is intended to briefly expose the differences
23 between a standard \LaTeX\ article and a paper based upon the
24 \texttt{cimento} class.
25 \end{abstract}

26

27 \section{Description}
28 This is a very short sample paper distributed with the class
29 \texttt{cimento}.
30 It is just a collection of examples about the syntax of commands
31 which behave in a different way from the standard \LaTeX
32 and/or new commands not defined in \LaTeX.

33

34 This sample is not meant to provide documentation for the class;
35 to obtain the class `manual' typeset the main distribution file,
36 \texttt{cimento.dtx}.

37

38 You can also use this file as a template for your own paper:
39 copy it to another filename and then modify as needed.

40

41 \section{Examples}
42 \subsection{Tables}
43 Tables~\ref{tab:pricesI}, \ref{tab:pricesII} and~\ref{tab:pricesIII}
44 inserted at this point.

The tables below show the usage of phantoms to align numbers. Also notice the absence of vertical and double horizontal lines.

45

46 \begin{table}
47 \caption{Prices of important items.}
48 \label{tab:pricesI}
49 \begin{tabular}{rcl}
50  \hline
51  Ice-cream      & 1500  & lire    \\
52  More ice-cream & 15000 & lire    \\
53  Crocodile      & 1500  & dollars \\
54  \hline
55  Phone call     & .25   & dollars \\
56  X-Men          & 1.25  & dollars \\
57  Dollar         & 1     & dollars \\
58  \hline
59 \end{tabular}
60 \end{table}

61

62 \begin{table}
63 \caption{Same thing, but formatted using \texttt{!} and \texttt{?}.}
64 \label{tab:pricesII}
65 \begin{tabular}{rcl}
66  \hline
67  Ice-cream      & !1500   & lire    \\
68  More ice-cream & 15000   & lire    \\
69  Crocodile      & !1500   & dollars \\
70  \hline
71  Phone call     & !.25    & dollars \\
72  X-Men          & 1.25    & dollars \\
73  Dollar         & 1?{.}!! & dollars \\
74  \hline
75 \end{tabular}
76 \end{table}

77

78 \begin{table}
79 \caption{Again, this time with \texttt{narrowtabular}}
80 \label{tab:pricesIII}
81 \begin{narrowtabular}{2cm}{rcl}
82  \hline
83  Ice-cream      & !1500   & lire    \\
84  More ice-cream & 15000   & lire    \\
85  Crocodile      & !1500   & dollars \\
86  \hline
87  Phone call     & !.25    & dollars \\
88  X-Men          & 1.25    & dollars \\
89  Dollar         & 1?{.}!! & dollars \\
90  \hline
91 \end{narrowtabular}
92 \end{table}

93

There is also an example of lettered equations. Please notice the usage of roman characters.

 95 \subsection{Mathematics}
 96 Here is a lettered array~(\ref{e.all}), with eqs.~(\ref{e.house})
 97 and~(\ref{e.phi}):
 98 \begin{eqnletter}
 99 \label{e.all}
100 \drm x_\sy{F} & = & 1.2\cdot10^3\un{cm}, \qquad
101  \tx{where\ } \sy{F} = \tx{Fermi}    \label{e.house}\\
102 \phi_i        & = & i\pi                                \label{e.phi}
103 \end{eqnletter}

104

105 \subsection{Citations}
106 We're almost done, just some citations~\cite{ref:apo}
107 and we will be over~\cite{ref:pul,ref:bra}.

108

The \appendix command changes the section headings as in standard L^{A}TEX.

109

110 \appendix

111

112 \section{}
113 Let us go then, you and I\ldots

114

115 \acknowledgments
116 This work was produced, supported and perpetrated by S. Frabetti under
117 the auspices of the Italian Physical Society.
118 Grazie to M. Missiroli for the valuable collaboration.

119

Now comes the bibliography. The macros \BY, \IN etc. are intended to help you match the Nuovo Cimento standards. Please note that the greatest part of the time spent in adapting an average L^{A}TEX document (including REVTEX and the like) goes in rearranging the references.

120 \begin{thebibliography}{0}
121 \bibitem{ref:apo} \BY{Boccaccio~G. \atque de~Cam\~oes~L.}
122 \IN{Phys. Rev. A}{13}{1999}{12};
123 \SAME{69}{999}{1666}.
124 \bibitem{ref:pul} \BY{Pulci~L.}
125 preprint INFN 8181.
126 \bibitem{ref:bra} \BY{Bragg~B.}
127 \TITLE{Tender comrade},
128 in \TITLE{Workers Playtime},
129  edited by \NAME{Tizio A. \atque Caio B.}
130  (Unexeditor, Bologna) 1997, pp.~1-10.
131 \end{thebibliography}

132

133 \end{document}

134 ⟨/sample⟩

References

[1] Leslie Lamport: L^{A}TEX: a document preparation system, 2nd ed.; Addison-Wesley, Reading, Massachussets, 1994 [2] Michel Goosens, Frank Mittelbach and Alexander Samarin: The L^{A}TEX Com-

panion; Addison-Wesley, Reading, Massachussets, 1994

[3] AA.VV., L^{A}TEX 2ε for authors, document distributed with L^{A}TEX 2ε under the

   file name usrguide.tex.